A flashlight on the bottom of a 3.95 m deep swimming pool sends a ray upward and at an angle so that the ray strikes the surface of the water 1.65 m from the point directly above the flashlight. What angle (in air) does the emerging ray make with the water's surface?
Answer #1

h-9.95m tono 145 0= 22-61 sin四二 0.985 - Now use Snell駕 law at the Watr _air ラ sin γ 0.5133 e= 90-2 90-30 , 81=53.13 Ans.